# Jasper and Stella: distillation of SOTA embedding models

### v1

@@ 3.1 Implementation details
Our model (jasper_en_vision_language_v1) is initialized from stella_en_1.5B_v5 and google/siglip-so400m-patch14-384Alabdulmohsin et al. (2024)Zhai et al. (2023). stella_en_1.5B_v5 and NV-Embed-v2 is our teacher models. The total number of parameters in the model is 1.9B (stella 1.5B parameters and siglip 400M parameters).
In all four stages, the model is trained with a maximum input length of 512 tokens, mixed precision training (BF16), DeepSpeed ZERO-stage-2 and AdamW optimizer.
In stage1 training (distillation training), the batch size is 128, the learning rate is 1e-04, trained with 8 RTX A6000, the checkpoint of step-4000 is the final model.
In stage2 training (distillation training), the batch size is 128, the learning rate is 8e-05, trained with 8 RTX A6000, the checkpoint of step-7000 is the final model.
In stage3 training (dimension reduction training), the batch size is 128, the learning rate is 7e-05, trained with 8 RTX A6000, the checkpoint of step-2200 is the final model.
In stage4 training (multimodal training), the batch size is 90, the learning rate is 1e-04, trained with 8 RTX A6000, the checkpoint of step-3500 is the final model.
@@ 3.2 Datasets
In stage1, stage2 and stage3, we use fineweb-eduLozhkov et al. (2024) as our main training dataset, which accounts for 80% of the full data. The remaining 20% of the data comes from sentence-transformers/embedding-training-data [^5] https://huggingface.co/datasets/sentence-transformers/embedding-training-data. The reason we choose sentence-transformers/embedding-training-data is that most fineweb-edu data is about passage; except for passage, we also need questions to enhance data diversity.
For the documents of data, we also do the following actions:
@@
    We randomly select 30% of the documents and cut them into short text (consisting of 1-10 sentences).
@@
    We randomly select 0.08% of the text and shuffle their words.
The total amount of data is 8 million.
In stage4, we use the caption data of BAAI/Infinity-MMGu et al. (2024) as our training data.
@@ 3.3 Results
We evaluate our model on the MTEB leaderboard and get an average score of 72.02 which is best in the models that have fewer than 2B parameters. The detailed results can be found in Table 1.
Model Average(56 datasets) Classification Clustering PairClassification Reranking Retrieval STS Summarization
NV-Embed-v2 72.31 90.37 58.46 88.67 60.65 62.65 84.31 30.7
jasper(our model) 72.02 88.49 58.04 88.07 60.91 63.12 84.67 31.42
bge-en-icl 71.67 88.95 57.89 88.14 59.86 62.16 84.24 30.77
stella_en_1.5B_v5 71.19 87.63 57.69 88.07 61.21 61.01 84.51 31.49
Table 1: MTEB Results
@@ 4 Discussion
In this section, we will discuss some interesting findings and possible improvements that have not yet been validated.
4.1 Robustness of instruction-based embedding models
The instruction-based embedding models mean that you should add an instruction to query or passage when encoding texts. Jasper model is also instruction-based. Currently, many excellent text embedding models use an instruction to prompt the model to get better embeddings. Just like usage of LLM, different tasks or scenarios use different instructions; this is both logical and intuitive. Hence, the ability to understand instructions is crucial to these text embedding models. In this subsection, we do a simple experiment on Jasper model to show the impact of different prompts. Specifically, we do a MTEB evaluation on some short evaluation-time tasks using similar instructions (generated by GPT-4o). Table 2 shows all original and modified instructions, and Table 3 shows the evaluation result. We think this evaluation result shows that Jasper model is robust and can correctly understand different instructions.
@@ 4.1 Robustness of instruction-based embedding models
Table 2: Original instructions and their synonyms
TaskType TaskName original_score score_with_modified_instructions
@@
Table 3: MTEB Results on different instructions
4.2 The Inconsistency of MSMARCO Scores
After releasing the Jasper model, an enthusiastic user (user name is raghavlite, https://huggingface.co/raghavlite) points out that the NDCG/MRR score is perfect and the MAP score is very low. Jasper model is distilled from stella_en_1.5B_v5 and NV-Embed-v2 and their MSMARCO score do not have this appearance. As of now, we still haven’t been able to figure out what happened. Anyway, we report this phenomenon in the hope that it will help some people. The details and updates can be found in https://huggingface.co/infgrad/jasper_en_vision_language_v1/discussions/3 .
@@ 4.3 Possible Improvements for Vision Encoding
Because of time and resource constraints, we were only able to give the jasper model a basic image encoding capability. In our initial conception, Stage4 is a basic visual language alignment training, while Stage5 (i.e. final stage, but we do not have this training) is the process of contrastive learning by using dataset of VQA. Besides, we found our loss in stage4 is oscillatory. All in all, there is a lot of room for improvement in the multimodal training stages.
@@ 5 Conclusion
In this paper, we introduce training details of Stella and Jasper models. The approaches of distillation and reducing dimensions can enable small models to achieve better results.

### v2

@@ 3.1 Implementation details
Our model named Jasper is initialized from stella_en_1.5B_v5 and google/siglip-so400m-patch14-384 Zhai et al. (2023); Alabdulmohsin et al. (2024). stella_en_1.5B_v5 and NV-Embed-v2 are our teacher models. The total number of parameters in our Jasper model is 1.9B (stella 1.5B parameters and siglip 400M parameters). For hyperparameters, we set $\lambda_{1}$ = 10, $\lambda_{2}$ = 200, $\lambda_{3}$ = 20, margin = 0.015.
In all four stages, the model is trained using 8 × RTX A6000 GPUs, with a maximum input length of 512 tokens, mixed precision training (BF16), DeepSpeed ZERO-stage-2, and the AdamW optimizer. During stage 1 (distillation training), the batch size is set to 128, the learning rate is 1e-4 per step, and the model checkpoint at step 4000 is selected as the final model. In the case of stage 2 (also distillation training), the batch size remains 128, the learning rate drops to 8e-5 per step, and the final model is the checkpoint at step 7000. For stage 3 (dimension reduction training), the batch size is again 128, the learning rate is adjusted to 7e-5 per step, and the checkpoint at step 2200 serves as the final model. Lastly, in stage 4 (multimodal training), the batch size is reduced to 90, the learning rate returns to 1e-4 per step, and the final model is chosen from the checkpoint at step 3500.
@@ 3.2 Datasets
In stage 1, stage 2 and stage 3, we use fineweb-edu Lozhkov et al. (2024) as our main text training dataset, which accounts for 80% of the full text data. The remaining 20% of the text data comes from sentence-transformers/embedding-training-data [^9] https://huggingface.co/datasets/sentence-transformers/embedding-training-data. The reason we choose the sentence-transformers/embedding-training-data is that the majority of the fineweb-edu data consists of passages. However, in addition to passages, we also require questions to enhance the diversity of our training data. The total amount of text training data is 8 million.
For the documents in our dataset, we perform the following actions:
@@
    We randomly select 30% of the documents and divide them into short texts, each consisting of 1 to 10 sentences.
@@
    We randomly select 0.08% of the text and shuffle the words within it.
In stage 4, we use the caption data of BAAI/Infinity-MM Gu et al. (2024) as our vision training data.
@@ 3.3 Results
We evaluate the proposed Jasper and Stella models on the full MTEB benchmark, which encompasses 15 retrieval datasets, 4 reranking datasets, 12 classification datasets, 11 clustering datasets, 3 pair classification datasets, 10 semantic textual similarity datasets, and 1 summarization dataset.
Table 1 presents the average score of our Jasper model across the overall performance and seven subcategory tasks of the METB benchmark. We compare our model with other frontier models on the MTEB leaderboard, as well as those with fewer than 2B parameters. Experimental results demonstrate that our Jasper model significantly outperforms other models with fewer than 2B parameters. Furthermore, despite having only 2B parameters, our model produces results that are comparable to those of models with 7B parameters.
@@ 4 Discussion
4.1 Instruction Robustness
Instruction-based embedding models require an instruction to be prepended to a query or passage during text encoding. Currently, many state-of-the-art text embedding models use instructions to prompt the model and obtain better embeddings. Similar to the usage of large language models Zhao et al. (2024b), different tasks necessitate different instructions, which is both logical and intuitive. Therefore, the ability to understand instructions is crucial for these text embedding models.
Jasper is also an instruction-based embedding model. To demonstrate the impact of different prompts on the Jasper model, we conducted a simple experiment. Specifically, we evaluated Jasper on some short evaluation tasks using similar instructions generated by GPT-4o. Table 2 lists all the original and modified instructions. Based on the results shown in Table 3, we conclude that our Jasper model is robust to instructions and can accurately understand different instructions.
@@ 4.1 Instruction Robustness
Table 2: Original instructions and corresponding synonyms.
Task Type Task Name Original Score Score with Modified Instructions
@@
Table 3: MTEB Results on different instructions.
@@ 4.2 Possible Improvements for Vision Encoding
Due to time and resource constraints, we were only able to equip the Jasper model with a basic image encoding capability. Initially, stage 4 was envisioned as a fundamental visual-language alignment training phase, with a potential stage 5 involving contrastive learning utilizing a Visual Question Answering (VQA) dataset. Additionally, we observed oscillatory behavior in our loss function during stage 4. Overall, there is considerable room for enhancement in the multimodal training.
@@ 5 Conclusion
In this paper, we present the distillation-based training procedure for the Jasper model. We have designed three loss functions to distill multiple large teacher embedding models into a student embedding model from diverse perspectives. Subsequently, we utilized a MRL-based training method to reduce the vector dimensionality of the student model. Experimental results on the MTEB demonstrate that our Jasper model achieves state-of-the-art performance at the 2B parameter scale and exhibits comparable results to other top-ranked embedding models with 7B parameters. Future work will further explore the alignment between multiple modalities.